[Brahms-l] Shift report 20060616 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 16 2006 - 00:12:14 EDT
Shifter leader was PS
Shifters were BB


1. no beam. tried two times to restore the beam but failed.
2. calibrated T3 and T4 for the whole run06 set. Have to work on T5 yet.
3. none.
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 16 00:14:20 2006

This archive was generated by hypermail 2.1.8 : Fri Jun 16 2006 - 00:14:34 EDT